OpenStudy (anonymous):

Hello, if arcos24/25, so what is result of cos? thanks, and how calculate that.

OpenStudy (lalaly):

arc cos is to find the angle, if u need the cos its cos(24/25) on the calculator,

OpenStudy (anonymous):

thanks

OpenStudy (aravindg):

nikvist you help me??

OpenStudy (anonymous):

Do you mean this?\[\arccos(24/25) = \theta\]\[\cos \theta = 24/25\]